OpenStudy (tukitw):

A cyclist accelerates steadily from rest on a 500 m downhill stretch of road, reaching a speed of 22.0 m s^-1 before free-wheeling to a halt 192 s after starting. Calculate how long the cyclist takes to reach top speed.

OpenStudy (compphysgeek):

do you have any information about the slope of the road after the first 500m?

OpenStudy (tukitw):

No

OpenStudy (compphysgeek):

If that's the only information that's given I have to assume that \(v = 22 ms^{-1}\) the maximum speed is. the acceleration was constant over \(s = 500 m\). The two equations we know for constant acceleration \(a\) is \[ v = a t \\ s = \frac{1}{2}a t^2\] we want to know the time it took to reach the maximum speed and we don't know the acceleration. So we solve the first equation for a and enter a into the second equation.\[ a = \frac{v}{t} \\ s = \frac{v}{2t}t^2 = \frac{vt}{2}\] We solve the equation for t \[t = \frac{2s}{v} \] Enter the known values into the final equation and you get the time it takes to accelerate to \(22 ms^{-1}\) over 500 m. Without any other information there is nothing else to calculate here. If the slope of the road is downwards after the 500 m the cyclist would still be accelerating and he would never stop. If the slope was horizontal or upwards the maximum speed would be 22 m/s. Dunno what the information of 192 seconds is there.

OpenStudy (tukitw):

Thank you, it's the correct answer. I was confused about the free-wheeling part.